OpenStudy (anonymous):

12x-8-7x+2=4 whats x and how did you find it?

OpenStudy (anonymous):

simplify to 5x-6=4 then simplify to 5x=10 x=10

OpenStudy (anonymous):

combine like terms

OpenStudy (katrinakaif):

Add like terms first: 12x - 7x - 8 + 2 = 4 4x - 6 = 4 Bring the -6 to the other side by adding it 4x = 4 + 6 4x = 10 Divide by 4 x = 10/4 Then simplify = 5/2

OpenStudy (anonymous):

12x-8-7x+2=4 5x-6=4 5x=4+6 5x=10 x=10/5 x =2

OpenStudy (katrinakaif):

AH sorry Abhishek is right..Substraction was wrong..Sorry!!!

OpenStudy (anonymous):

ohh, my bad, i agree, dont use my answer

OpenStudy (anonymous):

no problem kat:))))

OpenStudy (katrinakaif):

=) Thanks guys =)